OpenStudy (anonymous):

Identify the slope and the y-intercept of this equation; 6x-5y=-20

OpenStudy (anonymous):

6x-5y=-20 move x to the other side -5y=-6x-20 divide by -5 y=(6/5)x+4 slope is 6/5 y-intercept is 4

OpenStudy (solomonzelman):

y=mx+b m- slope b- y-intercept when ever you have this type of question, try to rearrange the equation, to make it look like y=mx+b or set it as y=.....

OpenStudy (anonymous):

ugh im confused

OpenStudy (solomonzelman):

I'll give you a different example, ready?

OpenStudy (anonymous):

ohh i got it noww

OpenStudy (solomonzelman):

Oh, OK, \(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\sf\color{blue}{~~~~Welcome~to~OpenStudy!~\ddot\smile}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (solomonzelman):

I'm here for about 40 days....